The Film Company uses a job-order costing system. The following information pertains to the company's activities for the month of December:

On December 1, the company had two jobs in work in process as follows: Job 6 $20,000 and Job 8 $14,000

On December 1, the company had job in Finished Goods as follows: Job 10 $50,000

Jobs 12 and 13 were started during the month.

Materials were purchased on account in December for $54,000

Materials were requisitioned for use in production. The direct materials were distributed as follows:

Job #6 $6,000 Job #8 $8,000 Job #12 $23,000 Job #13 $7,000

Direct labor was incurred in December. The direct labor was distributed as follows:

Job #6 $30,000 Job #8 $20,000 Job #12 $35,000 Job #13 $10,000

Overhead is applied using camera hours. The budgeted overhead for the coming year is $1,000,000

Estimated camera hours is the cost driver and is 250,000 camera hours

The following is the distribution of Camera hours (CH) for December:

Job #6 3,000 CH Job #8 4,000 CH Job #12 11,000 CH Job #13 2,000 CH

Actual overhead for December was incurred for various indirect costs on account.  

Use a generic miscellaneous payable account in your journal entry. $75,000

Job Nos. 8 and 12 were completed during the month.

Job Nos.10 and 12 were sold to customers on account during December at 125% of cost

Any over-or underapplied overhead is closed to Cost of Goods Sold.

Question

What is the ending balance in Work in Process on December 31. Give the applicable Job number(s), dollar amount for each job, and total ending Work in Process.

What is the ending balance in finished goods on December 31. Give the applicable Job number(s), dollar amount for each job, and total ending Finished Goods.

Calculate the cost of goods sold for the month ending December 31. Give the applicable Job number(s), dollar amount for each job, and total for Cost of Goods Sold.
Answer #1
Pre-determined overhead rate = Estimated Overhead/Estimated Camera hours
Pre-determined overhead rate =1000000/250000 = $4 per camera hour
JOB DETAILS:
Job 6 8 10 (FG OB) 12 13 Total Cost Selling price
OB 20000 14000 34000
D.MATERIAL 6000 8000 23000 7000 44000
D. LABOR 30000 20000 35000 10000 95000
OVERHEAD 3000*4=12000 16000                       44000 8000 80000
Total 68000 58000 50000 102000 25000 253000
COMPLETED FG 58000 58000
WIP WIP 68000 WIP 25000 93000
SOLD SOLD50000 SOLD102000 152000 152000*1.25=190000
Req 1) Ending Work in Process on Dec 31 =
Job 6 Job 13 Total
WIP 68000 25000 93000
Req 2) Ending FG on Dec 31 =
Job 8
Completed 58000
Req 3) COGS for the month Dec = Unadjusted COGS - Over-applied overhead
152000 - (80000 - 75000) = $147000